OpenStudy (anonymous):

the length of a rectangle is twice the width. the area is 648 yd^2 find the length and width the length of a rectangle is twice the width. the area is 648 yd^2 find the length and width @Mathematics

OpenStudy (anonymous):

648 = 2x(x) (l=2x, w = x) 648 = 2x^2 x^2= 324 x=324^1/2 x=18

OpenStudy (anonymous):

length = 36, width = 18, 36*18=648.